OpenStudy (anonymous):

An experiment consists of drawing a marble out of a jar. The jar has 10 orange marbles, 13 red marbles, 25 green marbles, and 32 blue marbles. Place the events in order from the least likely to occur to the most likely to occur. drawing a red or green marble drawing a yellow marble drawing a blue marble drawing any marble drawing an orange marble

OpenStudy (campbell_st):

order the colours based on the number of marbles. The more marbles of a colour, the more likely it is of being drawn

OpenStudy (texaschic101):

total marbles : 10 + 13 + 25 + 32 = 80 marbles chance of drawing an orange marble is : 10/80 which reduces to 1/8 or 12.5% chance of drawing a yellow arble : 0/80 because there is no yellow marbles -- 0% chance of drawing a blue marble : 32/80 which reduces to 2/5 or 40% chance of drawing any marble : 80/80 or 100% chance of drawing red marble : 13/80 = 16.25% chance of drawing a green marble : 25/80 which reduces to 5/16 = 31.25% chance of drawing red and green : I am not sure about this one...I do not know whether you add them or multiply them....13/80 (+ or x -- not sure)25/80